Feb 13, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and SeeVogh session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or www.readytalk.com (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Provide short summary of constants and their use in determination of the sampling fraction (David)
  2. Prepare a plan for cosmic-ray testing including external counters (Zisis)

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Monitoring
  4. FCAL
  5. BCAL
  6. Reconstruction and Simulation
  7. Regina
  8. TDR update
  9. Calibration
  10. Any other business


Attendees: Elton, John, Mark (JLab); George, Pavlos, Stratos (Athens); Shaun, Andrei (UofR); Sean (NW); Matt (IU)

  1. Announcements
  2. Action Items
    • No update
  3. Monitoring
    • Completed cabling and circuitry tested for FCAL LED monitoring system and found switched pin connections between distribution board and LED control board. Fuse blown on one FCAL control board, which can probably be fixed easily on site. We also have one spare control board
    • Four cables have been modified to account for the switched pin connections and LED system is operational.
    • Same problem exists for the BCAL monitoring system, where cable modifications are more difficult. The distribution boards will be modified instead. This will take a couple of weeks.
    • Orlando is working on the programing of the pulser system and scripts for various sequences can be written. For now the script will allow selection of one pulse line and its frequency.
  4. FCAL
    1. Interlock
      • All control system lights visible to pmts, so operation is now in total darkness (including crash buttons). LEDs in dark room are taped black during operation.
      • Currently, entry into the dark room only when HV is off. This limits some diagnostics, but essentially all operations are unaffected.
    2. Checkout
      • All pmt HV set to 1700 for diagnostics
      • Base check: 10 bases could not be programed (replaced); 5 had to be reprogramed (fixed); 5 have problems with setting the correct HV (should be replaced today).
      • Scripts to set individual HV settings on bases have been developed.
    3. DAQ
      • Only working trigger at the moment is the internal trigger
      • Sasha is working (with DAQ group) on getting external triggers to work - CODA problem
      • Pedestal studies have been performed for different configuration of pmts. Pedestal width is 1.3 counts (1.2 when bases are off), For details see FCAL DAQ Meetings
      • Mark: Could use random trigger (like FDC) if external trigger does not get fixed soon.
      • Dave has JANA working to parse and deliver data for analysis
    4. FCAL mini setup (Matt)
      • The FCAL prototype (used in beam test) is setup at IU, pointing up to the sky to investigate response to minimum ionizing
      • Graduate student on the project
      • Connected to Flash ADC for analysis
    5. Switching Noise
      • Switching noise from the base is visible on the pmt outputs.
      • Noise pulses visible every 2 micro seconds, corresponding to pulses internal to the Cockcroft-Walton base.
      • Macro duty cycle of these structures (expected to be a couple of KHz) is difficult to see
      • Effect on data seems to be small, considering minimal impact on pedestal width
  5. BCAL (Shawn)
  6. Reconstruction and Simulation
  7. Regina
    • Media:timing.pdf
    • Comparison of timing extracted from the Flash with that from the F1 TDCs was shown in attached plots.
  8. TDR update
  9. Calibration
  10. Any other business